While playing tag, Bob is being chased. He thinks he will get away faster by jumping on a stationary cart. Bob has a mass of 75kg and the cart has a mass of 25kg. If Bob’s speed is 4.0 m/s when he jumps on the cart, what is the speed of the cart after Bob jumps on?
Answer #1

Apply conservation g momentum, muitma U2 =/m, tma) v 7514) +0 = (75+ as) V= 3 m/s